OpenStudy (anonymous):

Given a wavelength of 121,270 millimeters, what is the wavelength in kilometers?

OpenStudy (anonymous):

1km=1000m 1m=100cm 1cm=10mm \[\frac{ 121,270 }{ 10 }cm=12127cm=\frac{ 12127 }{100 }m=121.27m=\frac{ 121.27 }{1000}km\]=.12127 km.

OpenStudy (anonymous):

thank you so much